OpenStudy (anonymous):

What is the solution set of |2x - 3| - 10 = -1? What is the solution set of |x| > 12?

OpenStudy (linknissan):

|2x-3|=9 2x-3=+or- 9

OpenStudy (linknissan):

2. x>+or-12

OpenStudy (anonymous):

1. x=6,-3 2. x>12 or x<-12

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

:)